¢15,000

Taxes paid to government 

¢5,000

Revenue received from sale of apples

¢35,000

Apples sold to the public

¢10,000

Apples sold to Kojo Juice Incorporation

¢25,000

Transactions of Kojo Juice Incorporation

Wages paid to workers of Kojo Juice Incorporation 

¢10,000

Taxes paid to government

¢2,000

Apples bought from Holi Apple Incorporation

¢25,000

Revenue received from sale of apple juice

¢40,000

The fundamental identity of national income accounting states that the same value of total 

economic activity is obtained whether activity is measured by the production of final goods and 

services (product approach), the amount of income generated by the economic activity (income 

approach), or the expenditure on final goods and services (expenditure approach). Using the data 

in the table above, show that the product approach, income approach and expenditure approach 

yield the same value of total economic activity (GDP).


1
Expert's answer
2021-06-23T12:54:41-0400

Wages to workers to pick apples by Appleinc "= \$15000" per year

Apples sold "= 25000+ 10000 = 35000" apples

Apple's profit before taxes "= 35000-15000 = \$20000"

Tax amount paid by Applelnc's"= \$5000"

After-tax profit of Appleinc"= \$15000"

Juicelnc buys 25000 apples from Apple and wages of $10000 are paid to workers in order to process the apples into apple juice.

It sells the apple juice for 40000, Juiceinc's Profit before taxes is"\$5000 = (40000 - 25000-10000)."

Taxes paid by Juiceinc "= \$2000"

After-tax profit of Juicelnc "= 5000 - 2000 = \$3000."


a) The product approach:

This approach is based on the concept of value that is added to the existing product. This could be calculated by subtracting the input value (purchased from the market) from the output. This computes the aggregates level of value added by all the producers within the economy.

This method also avoids the double-counting of economic activity.

In the given question:

The output value of Applelnc is 35000 and that of Juicelnc is 40000.

Now we will sum the value-added"\$(35000 + (40000 - 25000)) = \$(35000+ 15000) = \$50000"


b) The Income Approach:

This approach is based on the concept of adding all the incomes of producers, such as wagers received by workers and the profit earned by the owners. as well as adding the revenues of the government which is essentially is called the tax.

"G.D.P. = \$(15000+ 15000 + 10000 + 3000 + 7000) =\$50000."


c) The expenditure approach

This approach is based on the spending pattern, which the expenditure incurred by consumption, investment (purchasing of raw material), government expenditure, and next export.

"G.D.P. = \$(10000 + 25000 + 15000 )= \$50000"

Answer d)

As the answer of parts a), b) and c) is equal, All the three approaches give the same value of G.D.P. because it essentially gives the market value of final goods and services, from different angles.